Please help me!
Evaluate the function for the given value of x
f(x)=3x+12;-5
g(x)=2.25x;100


Answer by stanbon(75887)

Evaluate the function for the given value of x
f(x)=3x+12;-5
f(-5) = 3(-5)+12 = -15+12 = -3
---------------------------------------
g(x)=2.25x;100
g(100) = 2.25*100 = 225
============================
Cheers,
Stan H.
============================